OpenStudy (anonymous):

Solve. 13-0.3(j-9)=0.9(11j-9)

OpenStudy (anonymous):

ick decimals

OpenStudy (anonymous):

multiply both sides by 10 to get rid of them stsart with \[130-3(j-9)=9(11j+9)\] and solve that one

OpenStudy (anonymous):

\[130-3j+27=99j+81\] etc